OpenStudy (anonymous):

write a unit rate for the situation. round to the nearest hundredth if necessary traveling 216 miles in 5 hours A( 54 miles/hours B( 39.27 miles/hours C( 43.2 miles/hours D( 0.4 miles/hours

zepdrix (zepdrix):

\[\large\rm \frac{216miles}{5hours}\qquad=\qquad\frac{216}{5}~\frac{miles}{hour}\]Simply divide to find a unit rate :)

jhonyy9 (jhonyy9):

216 ----- 5 h x ------ 1 h ------------- x = 216/5 = ?